five times the sum of two consecutive odd integers is twelve times the larger of the two. find the two odd integers.
Answer by peter_mccarter(5)

Five times the sum of two consecutive odd integers is twelve times the larger of the two. find the two odd integers.
Remember that two odd consecutive integers are going to be two numbers apart 1,3,5,7 ect.
1st number = X and the 2st number = x+2
5(x+x+2)=12(x+2) Now do the work inside the parentheses 5(2x+2)=12(x+2)
Now distribute the 5 and the 12 in to the parentheses 10x+10=12x+24
Now subtract 10x and 24 from both sides -14=2x now to isolate the x divide by 2
-14 ÷2= -7 , 2x÷2=x now plug into the original equation
1st number =x
2nd number =x+2 or -7+2=-5
X=-7
X+2= -5